# Improved quantum data analysis

### 12 Dec 2023 v3

@@ 5.2 Hypothesis Selection
In this section we establish our quantum Hypothesis Selection result, Theorem 1.5. This theorem effectively has three different bounds, and we prove them via Propositions 5.3, 5.5 and 5.6.
@@
One way to solve the quantum Hypothesis Selection problem is to simply use Shadow Tomography as a black box. Given parameters $0<\epsilon,\delta<\frac{1}{2}$ for the former problem, we can run Shadow Tomography with parameters $\epsilon/2$, $\delta$, and the $\binom{m}{2}$ quantum events $(A_{ij}:i<j)$. Then except with probability at most $\delta$, we obtain values $\widehat{\mu}_{ij}$ with $\lvert\operatorname*{\mathbf{E}}_{\rho}[A_{ij}]-\widehat{\mu}_{ij}\rvert\leq\epsilon/2$ for all $i,j$. Now we can essentially use any classical Hypothesis Selection algorithm; e.g., the “minimum distance estimate” method of Yatracos [Yat85]. We select as our hypothesis $\sigma_{k}$, where $k=\operatorname*{argmin}_{\ell}\widehat{\Delta}_{\ell}$ is a minimizer of
@@
where the last inequality used the Holevo–Helstrom theorem again, and the Shadow Tomography guarantee. We now obtain the following result (with the proof being an almost verbatim repeat of the one in [DL01, Thm. 6.3]):
@@
and the result now follows from Equation 4. ∎
@@ Appendix A The quantum Threshold Decision problem
As mentioned at the end of Section 1.1.1, Aaronson [Aar20] showed that the decision version of quantum Threshold Search can be done with $n=O(\log(m)\log(1/\delta)/\epsilon^{2})$ copies, through the use of a theorem of Harrow, Lin, and Montanaro [HLM17, Cor. 11]. In Theorem A.2 below, we give a new version of the Harrow–Lin–Montanaro theorem, with a mild qualitative improvement. This improvement also lets us improve the quantum Threshold Decision copy complexity slightly, to $n=O(\log(m/\delta)/\epsilon^{2})$ (see Corollary A.4).
@@
$=\sqrt{\operatorname{tr}(\rho X)}\sqrt{\operatorname{tr}(\rho Y^{\dagger}XY)}.\qed$
@@
One can read out a similar result in the work of Harrow, Lin, and Montanaro [HLM17, Cor. 11], except with a lower bound of $\operatorname*{\mathbf{E}}_{\rho}[B]\geq.632(p_{\max}-\nu)^{2}$. Note that unlike our bound, their lower bound is never close to $1$, even when $p_{\max}$ is very close to $1$. It is this difference that leads to our slight improvement for the Threshold Decision problem. We speculate that the lower bound in our result can be sharpened further, to $(1-O(\sqrt{\nu}))p_{\max}$.
Proof.
@@
$p\leq\sqrt{\beta}+\sqrt{\delta}\sqrt{p}<\sqrt{\nu}+\frac{\delta+p}{2}\implies p\leq 2\sqrt{\nu}+\delta.\qed$
Given Theorem A.2, it’s easy to obtain the following quantum Threshold Decision algorithm, similar to [Aar20, Lem. 14]:
@@
This can be solved using just $n=O(\log(m/\delta)/\epsilon^{2})$ copies of $\rho$ and probability of error at most $\delta$. The algorithm can be implemented by a projector applied to $\rho^{\otimes n}$.)
@@
$\lvert\operatorname*{\mathbf{E}}_{\rho}[A_{i}]-\theta_{i}\rvert>\eta_{i}+\epsilon\implies\operatorname*{\mathbf{E}}_{\varrho}[A^{\prime}_{i}]\geq 1-\delta/2,\qquad\lvert\operatorname*{\mathbf{E}}_{\rho}[A_{i}]-\theta_{i}\rvert\leq\eta_{i}\implies\operatorname*{\mathbf{E}}_{\varrho}[A^{\prime}_{i}]\leq\delta^{3}/(16m).\qed$
a>

### 15 Mar 2024 v4

@@ 5.2 Hypothesis Selection
In this section we establish our quantum Hypothesis Selection result, Theorem 1.5. This theorem effectively has three different bounds, and we prove them via Propositions 5.5, 5.8 and 5.9.
@@
One way to solve the quantum Hypothesis Selection problem is to simply use Shadow Tomography as a black box. Given parameters $0<\epsilon,\delta<\frac{1}{2}$ for the former problem, we can run Shadow Tomography with parameters $\epsilon/2$, $\delta$, and the $\binom{m}{2}$ quantum events $(A_{ij}:i<j)$. Then except with probability at most $\delta$, we obtain values $\widehat{\mu}_{ij}$ with $\lvert\operatorname*{\mathbf{E}}_{\rho}[A_{ij}]-\widehat{\mu}_{ij}\rvert\leq\epsilon/2$ for all $i,j$. Now we can essentially use any classical Hypothesis Selection algorithm; e.g., the “minimum distance estimate” method of Yatracos [40]. We select as our hypothesis $\sigma_{k}$, where $k=\operatorname*{argmin}_{\ell}\widehat{\Delta}_{\ell}$ is a minimizer of
@@
where the last inequality used the Holevo–Helstrom theorem again, and the Shadow Tomography guarantee. We now obtain the following result (with the proof being an almost verbatim repeat of the one in [14, Thm. 6.3]):
@@
and the result now follows from Equation 4.
@@ Appendix A The quantum Threshold Decision problem
As mentioned at the end of Section 1.1.1, Aaronson [1] showed that the decision version of quantum Threshold Search can be done with $n=O(\log(m)\log(1/\delta)/\epsilon^{2})$ copies, through the use of a theorem of Harrow, Lin, and Montanaro [24, Cor. 11]. In Theorem A.3 below, we give a new version of the Harrow–Lin–Montanaro theorem, with a mild qualitative improvement. This improvement also lets us improve the quantum Threshold Decision copy complexity slightly, to $n=O(\log(m/\delta)/\epsilon^{2})$ (see Corollary A.6).
@@
$=\sqrt{\operatorname{tr}(\rho X)}\sqrt{\operatorname{tr}(\rho Y^{\dagger}XY)}.$
@@
One can read out a similar result in the work of Harrow, Lin, and Montanaro [24, Cor. 11], except with a lower bound of $\operatorname*{\mathbf{E}}_{\rho}[B]\geq.632(p_{\max}-\nu)^{2}$. Note that unlike our bound, their lower bound is never close to $1$, even when $p_{\max}$ is very close to $1$. It is this difference that leads to our slight improvement for the Threshold Decision problem. We speculate that the lower bound in our result can be sharpened further, to $(1-O(\sqrt{\nu}))p_{\max}$.
Proof A.5.
@@
$p\leq\sqrt{\beta}+\sqrt{\delta}\sqrt{p}<\sqrt{\nu}+\frac{\delta+p}{2}\implies p\leq 2\sqrt{\nu}+\delta.$
Given Theorem A.3, it’s easy to obtain the following quantum Threshold Decision algorithm, similar to [1, Lem. 14]:
@@
This can be solved using just $n=O(\log(m/\delta)/\epsilon^{2})$ copies of $\rho$ and probability of error at most $\delta$. The algorithm can be implemented by a projector applied to $\rho^{\otimes n}$.
@@
$\lvert\operatorname*{\mathbf{E}}_{\rho}[A_{i}]-\theta_{i}\rvert>\eta_{i}+\epsilon\implies\operatorname*{\mathbf{E}}_{\varrho}[A^{\prime}_{i}]\geq 1-\delta/2,\qquad\lvert\operatorname*{\mathbf{E}}_{\rho}[A_{i}]-\theta_{i}\rvert\leq\eta_{i}\implies\operatorname*{\mathbf{E}}_{\varrho}[A^{\prime}_{i}]\leq\delta^{3}/(16m).$